# DrugMCTS: a drug repurposing framework combining multi-agent, RAG and Monte Carlo Tree Search

### v2

@@ Abstract
Recent advances in large language models have demonstrated considerable potential in scientific domains such as drug discovery. However, their effectiveness remains constrained when reasoning extends beyond the knowledge acquired during pretraining. Conventional approaches, such as fine-tuning or retrieval-augmented generation, face limitations in either imposing high computational overhead or failing to fully exploit structured scientific data. To overcome these challenges, we propose DrugMCTS, a novel framework that synergistically integrates RAG, multi-agent collaboration, and Monte Carlo Tree Search for drug repurposing. The framework employs five specialized agents tasked with retrieving and analyzing molecular and protein information, thereby enabling structured and iterative reasoning. Without requiring domain-specific fine-tuning, DrugMCTS empowers Qwen2.5-7B-Instruct to outperform Deepseek-R1 by over 20%. Extensive experiments on the DrugBank and KIBA datasets demonstrate that DrugMCTS achieves substantially higher recall and robustness compared to both general-purpose LLMs and deep learning baselines. Our results highlight the importance of structured reasoning, agent-based collaboration, and feedback-driven search mechanisms in advancing LLM applications for drug discovery.
@@ 1 Introduction
As interest grows in applying general-purpose large models to scientific domains, various approaches have been proposed to address the aforementioned limitations. Among them, fine-tuning on domain-specific datasets has become a widely adopted paradigm [3] [4]. However, this approach comes with several notable drawbacks. First, fine-tuning demands substantial computational resources, and most current methods are tailored to specific domains. Extending these models to new domains typically requires additional training, which is computationally inefficient and often impractical for many users.
Moreover, scientific knowledge is inherently dynamic, constantly evolving with new discoveries. This necessitates continuous updates to the model’s internal knowledge base, which in turn introduces challenges such as increased training costs and the risk of catastrophic forgetting [5]. These issues significantly limit the scalability and practicality of fine-tuning-based approaches in real-world scientific applications.
To mitigate the limitations of fine-tuning, an alternative paradigm, retrieval-augmented generation (RAG), has gained increasing attention [6]. In this framework, external agents are employed to retrieve relevant information from scientific literature and databases, which is then used to augment the reasoning capabilities of LLMs [7]. This approach eliminates the need for fine-tuning and enables access to up-to-date knowledge, making it particularly appealing for fast-evolving domains like drug discovery.
However, significant challenges remain. As noted in prior work [38], Data in scientific domain can be broadly categorized into two types: scientific data and general-purpose data. Most existing RAG-based systems heavily rely on the latter due to its compatibility with general-purpose LLMs [8], while largely overlooking the value of structured, authoritative scientific data, such as molecular structures and protein sequences, which is often cleaner and more informative than generic text. For instance, in drug-target interaction prediction tasks, some approaches only provide knowledge graphs or textual descriptions as context, completely omitting structural information about proteins or molecules [10]. This omission undermines both the reliability of the model’s predictions and the interpretability of its decision-making process.
Some approaches incorporate domain-specific models during the reasoning phase to better understand scientific data [9]. While effective in certain contexts, this strategy inherits many of the limitations of the fine-tuning paradigm. Notably, when the input data deviates from the distribution seen during training, the performance of these models can degrade significantly. For example, deep learning models for drug-target interaction prediction may experience accuracy drops of over 20% when encountering previously unseen molecule-protein pairs [11].
On the other hand, general-purpose data sources often contain irrelevant or even erroneous information [12], necessitating preprocessing steps such as filtering and cleaning before use. However, LLMs may inadvertently discard useful content due to incomplete contextual understanding or internal biases. As highlighted in [1], drug discovery should ideally be an iterative process driven by feedback and refinement. Yet, most current approaches rely on single-step inference without mechanisms for error correction or knowledge updating, limiting their robustness and adaptability [36] [37].
To address the aforementioned limitations, we propose DrugMCTS, a novel drug discovery algorithm based on RAG, multi-agent collaboration, and Monte Carlo Tree Search (MCTS) [34]. Our system comprises five specialized agents:
  • Retrieval Agent: Identifies and gathers potentially relevant molecules.
  • Molecule-Analysis Agent: Evaluates the physicochemical and pharmacological properties of the query molecule.
  • Molecule-Selection Agent: Filters out molecules with limited therapeutic potential or low relevance.
  • Interaction-Analysis Agent: Interprets the molecular mechanisms underlying drug-target interactions.
  • Decision Agent: Integrates all available evidence and generates final recommendations.
With the guidance of MCTS, our lightweight Qwen7b-based model [40] achieves performance exceeding that of Deepseek-R1 [33] and GPT-4o-mini [32], surpassing it by approximately 20% on key benchmarks. We also conduct ablation studies demonstrating that removing any component of our pipeline results in a performance drop of 2–10%. Our key contributions include:
@@
    We introduce an end-to-end drug discovery framework that enables Qwen2.5-7B-Instruct to outperform much larger models Deepseek-R1. The method does not require any domain-specific fine-tuned models; instead, it leverages external knowledge to enhance reasoning and decision-making.
@@
    We propose a systematic workflow that integrates scientific data, hybrid scientific-general data, and general-purpose data, leveraging the strengths of each type. This pipeline not only improves the interpretability of model decisions but also establishes a standardized framework applicable beyond drug-target interaction tasks.
@@
    By incorporating a feedback mechanism through Monte Carlo Tree Search, our framework enables the model to iteratively refine its understanding, filter noisy data, and autonomously identify the most valuable information for decision-making.
@@ 2 Method
2.1 Overview
Our framework, DrugMCTS, as shown in Figure 1, takes a query molecule $M_{qm}$ as input and processes it through a series of specialized agents to output proteins that can potentially interact with $M_{qm}$. The process begins with the Retrieval Agent, which queries databases based on the molecular structure of the $M_{qm}$ to identify similar molecules, thereby generating a pool of candidate molecules $M_{cm}$. Next, the Molecule-Analysis Agent retrieves chemical properties of the query molecule via API calls and generates a detailed report $R_{qm}$. This report is then passed to the Molecule-Selection Agent, which filters the $M_{cm}$ based on the analysis, producing a refined list of reference molecules $M_{rm}$. The Interaction-Analysis Agent identifies proteins $P_{rp}$ that can interact with the reference molecules $M_{rm}$ by querying relevant databases. For each identified protein-molecule pair, the agent retrieves the corresponding binding pocket data $D_{bp}$ from the Protein Data Bank (PDB) [13] using API calls and python libraries. Additionally, it fetches literature descriptions $L_{rp}$ related to the proteins from public repositories such as PubMed. It then compiles all this information, $M_{rm}$, $P_{rp}$, $P_{bp}$ and $L_{rp}$, into a comprehensive interaction report $R_{ia}$ for each molecule-protein pair.
Finally, the Decision Agent integrates all available information, including the $M_{qm}$, $R_{qm}$, $M_{rm}$, $D_{bp}$, and $R_{ia}$, to select the most promising target protein.
To enhance decision-making, we employ MCTS during inference time scaling. Unlike the retrieval agent, other agents generate multiple answers in a single invocation. These answers are evaluated using the Upper Confidence Bound applied to Trees (UCT) algorithm (Equation 1[19] to select and expand optimal nodes until reaching the leaf node. We use a self-confidence scoring method for leaf nodes and backpropagate scores accordingly. The final selection of proteins is determined by majority voting, choosing the top N proteins that appear most frequently. This approach allows our model to autonomously select high-quality data without additional fine-tuning, effectively achieving data cleaning through feedback mechanisms.
@@ 2.1 Overview
Figure 1: Workflow of DrugMCTS
2.2 Data Processing Pipeline
In prior work [38], data in scientific domain has been categorized into two types: scientific data and general data. Scientific data includes highly specialized information such as molecular formulas and amino acid sequences, characterized by its authoritative nature, well-structured format, and cleanliness. These datasets are typically sourced from professional databases like the PDB [13], where data storage formats adhere to strict standards. However, LLMs often struggle to interpret such structured data directly due to its complexity, necessitating the use of domain-specific models like ESM2 [14] or ChemBERTa [15] for processing.
Conversely, general data encompasses more diverse and less standardized sources, such as research papers, which are easier for LLMs to understand but may contain noise and inaccuracies. Leveraging the strengths of both data types while mitigating their weaknesses is crucial for effective drug discovery.
Based on the original work, we propose a new data type called hybrid scientific-general data. Tools like RDKit [16] can accept scientific inputs, such as SMILES representations of molecules, and output structured yet textually described data, including chiral centers, functional groups, and Murcko scaffolds, packaged as Python dictionaries. Similarly, PLIP [17] methods can process PDB files to extract pocket information and return it in tabular form. These hybrid data outputs retain the authority and well-structured nature of scientific data while being more interpretable by LLMs due to their textual explanations.
Our workflow involves transforming both molecular and protein data through three stages: from scientific data to hybrid scientific-general data, and finally to general data. This structured pipeline ensures that our model benefits from the precision and reliability of scientific data while maintaining interpretability and ease of integration with LLMs. By systematically leveraging these different data types, we not only improve the accuracy and robustness of our predictions but also enhance the explainability of our model’s decision-making process.
2.3 Formulating Drug Discovery as a Tree Search Problem
Given that MCTS is a well-established algorithm widely adopted and thoroughly explained in various studies [18], we provide only an overview here. The MCTS algorithm consists of four main steps: selection, expansion, simulation, and backpropagation.
@@ 2.4 Action Space
$A_{1}$ Retrieval Action. Upon receiving the query molecule, the Retrieval agent queries databases to identify molecules structurally similar to the query molecule. We employ two similarity metrics: the Tanimoto coefficient [29] and the cosine similarity based on the last hidden state computed by ChemBERTa. For each metric, we retrieve the top-10 most similar molecules, merge the results, and remove duplicates to form the initial candidate molecules $M_{cm}$. The proteins can interact with $M_{cm}$ are also retrieved, forming candidate proteins $P_{cp}$. These molecules serve as inputs to the root node of the Monte Carlo tree search.
@@
$A_{2}$ Molecule Analysis Action. Understanding molecular properties is crucial for predicting molecular-target interactions. While general-purpose LLMs can perform some analysis using SMILES representations, their interpretations are often incomplete and prone to errors, especially when experimental data, such as hydrophobicity, is required.
@@
$R_{qm}=\text{MA Agent}(M_{qm},C_{q,s},C_{q,phy})$ (4)
$A_{3}$ Molecule Selection Action. Research has shown that the quality of retrieved information significantly impacts the accuracy of model-generated answers. Excessive irrelevant information can negatively affect the model’s performance. Although the retrieved molecules share structural similarities with the query molecule, they may not necessarily provide useful insights for drug discovery tasks. Thus, the Molecule Selection (MS) agent filters the molecule pool, based on structural similarity to the query molecule, pharmacophore integrity, and drug-like essentials, to generate reference molecules $M_{rm}$. The reference proteins $P_{rp}$ are obtained by selecting the proteins that can interact with $M_{rm}$ from $P_{cp}$. To ensure that these reference molecules are thoroughly characterized, this action also invokes the $A_{2}$ Molecule Analysis Action. Specifically, the MA agent is called to retrieve the structural and physicochemical properties ($C_{c,s}$ $C_{c,phy}$) of $M_{cm}$ but without generating reports.
$M_{rm}=\text{MS Agent}(M_{qm},R_{qm},M_{cm},C_{c,s},C_{c,phy})$ (5)
@@ 2.5 Reward Calculation
Molecular Docking Methods. Molecular docking methods, such as AutoDock Vina [23], are widely used but suffer from notable limitations. First, they require accurate three-dimensional (3D) structural information of the target proteins, which is often unavailable or unreliable for many biologically relevant molecules. Second, these methods are computationally intensive, limiting their scalability in large-scale drug discovery tasks. Given these limitations, there has been a growing interest in developing more efficient alternatives.
Deep Learning-Based Methods. Deep learning models have emerged as a promising solution due to their superior computational efficiency [39]. However, their performance is highly dependent on the training dataset. When predicting samples that deviate significantly from the training data, the accuracy of these models can drop by over 30%, limiting their applicability across diverse scenarios.
@@ 3.1 Datasets and Metrics
We utilized two datasets, DrugBank [31] and KIBA [30], which were processed to include a total of 788 entries from DrugBank and 626 entries from KIBA. Each entry consists of a molecule as input and its corresponding interacting proteins as output. The number of ground truth interactions per entry varies. To evaluate model performance, we used recall, defined as the ratio of correctly predicted proteins to the total number of ground truth proteins. The detailed information of data processing pipeline can be found in Appendix Section A.1
Recall=|{proteins predicted correctly}||{all ground truth $\text{Recall}=\frac{|\{\text{proteins predicted correctly}\}|}{|\{\text{all ground truth proteins}\}|}$ (11)
@@ 3.3 Baselines
General Models with RAG. (GM + RAG) Many existing studies on drug repurposing using LLMs either employ divergent methodological formulations or are not open-sourced, which complicates comparative evaluation. Our approach adopts the framework of the DrugReAlign [22] method, with modifications and enhancements tailored to our specific formulation. To ensure equitable comparison, we incorporate not only protein pocket information but also incorporate $C_{q,s}$ and $C_{q,phy}$ as additional features.. The models used in this group remained GPT-4o-mini and Deepseek-R1.
@@ 3.4 Results
The experimental results (Figure 1) indicate that general-purpose LLMs (GPT-4o-mini and Deepseek-R1) achieved relatively low recall scores of only 12.59%–16.19% on the DrugBank dataset when operating in a zero-shot setting. However, when incorporating molecular structural features and chemical properties via RAG-based prompting, model performance decreased, with GPT-4o-mini dropping to 15.19% and Deepseek-R1 significantly falling to 12.59%. This suggests that the inclusion of potentially irrelevant or misleading information through retrieval can negatively impact the reasoning capabilities of general-purpose LLMs.
@@
These results strongly underscore the superiority of dynamic decision-making mechanisms, such as those used in DrugMCTS, over traditional static prediction methods like deep learning models. They also highlight the limitations of general-purpose large models in zero-shot settings for drug discovery tasks, especially when retrieval-augmented prompting introduces noise or irrelevant context. This further emphasizes the importance of structured reasoning, domain-specific knowledge integration, and information filtering in such applications.
Furthermore, to demonstrate the interpretability and transparency of our approach, we provide a detailed case study in the Appendix Section A.2. It illustrates how the model predicts an interaction between Equol and the CXC chemokine receptor 3, including the complete step-by-step reasoning process during a specific MCTS rollout.

### v3

# DrugMCTS: a Drug Repositioning Framework Combining Multi-Agent, RAG and Monte Carlo Tree Search

@@ Abstract
Recent advances in large language models have demonstrated considerable potential in scientific domains such as drug repositioning. However, their effectiveness remains constrained when reasoning extends beyond the knowledge acquired during pretraining. Conventional approaches, such as fine-tuning or retrieval-augmented generation, face limitations in either imposing high computational overhead or failing to fully exploit structured scientific data. To overcome these challenges, we propose DrugMCTS, a novel framework that synergistically integrates RAG, multi-agent collaboration, and Monte Carlo Tree Search for drug repositioning. The framework employs five specialized agents tasked with retrieving and analyzing molecular and protein information, thereby enabling structured and iterative reasoning. Extensive experiments on the DrugBank and KIBA datasets demonstrate that DrugMCTS achieves substantially higher recall and robustness compared to both general-purpose LLMs and deep learning baselines. Our results highlight the importance of structured reasoning, agent-based collaboration, and feedback-driven search mechanisms in advancing LLM applications for drug repositioning. Contact: zeruiyang2-c@my.cityu.edu.hk
@@ 1 Introduction
As general-purpose large models gain traction in scientific domains, fine-tuning on domain-specific datasets has emerged as a common approach [36, 27]. However, this method faces several well-recognized limitations. Fine-tuning is computationally expensive and typically domain-specific, making it inefficient and often impractical for extending to new scientific areas. Furthermore, given the dynamic nature of scientific knowledge, maintaining up-to-date models requires continual retraining, which increases computational overhead and risks catastrophic forgetting [19]. These challenges hinder the scalability and long-term utility of fine-tuning in real-world scientific applications.
To address the drawbacks of fine-tuning, retrieval-augmented generation (RAG) has emerged as a promising alternative [35]. By leveraging external agents to retrieve relevant information from literature and databases, RAG enhances the reasoning capabilities of LLMs without requiring model updates [3]. This makes it particularly suitable for dynamic domains such as drug discovery.
Despite its advantages, RAG faces notable limitations. Scientific data can be broadly categorized into structured scientific data and unstructured general-purpose text [39]. Existing RAG systems often prioritize the latter due to better compatibility with LLMs [23], overlooking the richness of structured data such as molecular structures and protein sequences. For example, drug-target interaction tasks frequently exclude structural information, relying solely on knowledge graphs or text [14], which compromises both prediction reliability and interpretability.
Some methods incorporate domain-specific models to interpret scientific data [11, 17], but this reintroduces fine-tuning-related limitations. When inputs deviate from training distributions, performance can drop sharply—for instance, drug-target interaction models may lose over 20% accuracy on unseen molecule-protein pairs [31].
General-purpose data also presents challenges. It often contains noise or errors [10], and LLMs may inadvertently discard useful content due to limited contextual understanding. For instance, long-context models are known to suffer from the “lost in the middle” [16] phenomenon (Figure 1), where information located in the middle of a long input sequence is more likely to be forgotten or overlooked [16], leading to incomplete or inaccurate reasoning. Moreover, current approaches typically lack feedback loops, treating drug discovery as a one-shot task rather than an iterative process [33, 4, 6], which limits robustness and adaptability.
Figure 1: When the document relevant to the query appears in the middle of the retrieved document sequence, the model’s performance can even fall below that of the baseline closed-book setting [16].
To address the aforementioned limitations, we propose DrugMCTS, a novel drug repositioning algorithm based on RAG, multi-agent collaboration, and Monte Carlo Tree Search (MCTS) [2]. DrugMCTS is designed to overcome the challenges posed by computationally expensive fine-tuning, underutilization of structured scientific data, model brittleness in out-of-distribution scenarios, and the lack of iterative feedback mechanisms.
Our system comprises five specialized agents:
  • Retrieval Agent: Identifies and gathers structurally similar molecules from external scientific databases. While this is the first retrieval step in the pipeline, subsequent agents also incorporate retrieval mechanisms to augment the model’s input with task-relevant scientific knowledge. By leveraging external sources without modifying model parameters, these retrieval-based actions collectively mitigate the limitations of static knowledge and the high cost of fine-tuning.
  • Molecule-Analysis Agent: Receives a dictionary containing the physicochemical and structural properties of the query molecule and its candidates, obtained from scientific APIs (e.g., RDKit, PubChem). This agent then summarizes the input into a concise natural language description, making the information more accessible to LLMs.
  • Molecule-Selection Agent: Filters retrieved molecules based on pharmacophore integrity and drug-likeness, effectively eliminating low-relevance or noisy inputs that could compromise model performance. This filtering step also helps mitigate the “lost in the middle” phenomenon.
  • Interaction-Analysis Agent: Analyzes drug-target interaction mechanisms by extracting structured protein data (e.g., pocket features from PDB files via PLIP) and enriching it with natural language summaries from literature. This transformation enables general LLMs to reason over biological structures that they cannot natively parse.
  • Decision Agent: Integrates all upstream outputs—including molecular properties, interaction analyses, and supporting texts—and makes final predictions about likely protein targets. It serves as the final reasoning step informed by the entire multi-agent pipeline.
Moreover, the incorporation of MCTS enables the framework to operate with a feedback-driven mechanism, iteratively refining its decision path. This allows the system to explore multiple reasoning trajectories and prioritize more promising nodes based on reward signals, thereby improving both robustness and decision quality over time.
With the guidance of our framework, the lightweight Qwen2.5-7B-Instruct model [26] achieves performance surpassing Deepseek-R1 [7], GPT-4o-mini [9] and domain-specific deep learning models on benchmark tasks. Notably, this is accomplished without any domain-specific fine-tuning. We further conduct ablation studies demonstrating that removing any component of our pipeline leads to a performance drop of 1–10%, underscoring the necessity of each module. To facilitate future research and ensure reproducibility, the code and data processing pipeline for DrugMCTS will be released upon publication at: https://github.com/yaoge777/DrugMCTS
Our main contributions include:
@@
    We introduce an end-to-end drug repositioning framework that enables Qwen2.5-7B-Instruct to outperform much larger models such as Deepseek-R1, without requiring any fine-tuned domain-specific model. The method leverages external structured and textual knowledge sources to boost reasoning and decision-making capabilities.
@@
    By incorporating MCTS, our framework introduces a feedback mechanism that enables iterative refinement of decisions, autonomous filtering of noisy inputs, and robust identification of high-value evidence—addressing the lack of adaptability in current one-shot inference systems.
@@
    We propose a systematic data processing pipeline (Figure 2) that transitions from structured scientific data to hybrid scientific-general data, and finally to general-purpose language input. This hybrid representation exploits the strengths of each modality, enhances model interpretability, and offers a reusable workflow applicable beyond drug-target interaction tasks.
Figure 2: Data processing pipiline of DrugMCTS
@@ 2 Method
2.1 Overview of DrugMCTS
Our proposed framework, DrugMCTS, enhances molecular-protein interaction prediction through a structured multi-agent system guided by MCTS, as illustrated in Figure 3 and Appendix Algorithm 1. Given a query molecule $M_{qm}$, DrugMCTS coordinates a sequence of agents to retrieve relevant chemical and biological information, perform structured analysis, and identify promising protein targets.
Each agent contributes domain-specific reasoning: retrieving structurally similar molecules, analyzing molecular properties, filtering irrelevant candidates, and evaluating interaction potential through structural and textual evidence. Instead of relying on end-to-end model predictions, DrugMCTS incrementally builds reasoning paths, ensuring interpretability and robustness at every step.
Crucially, we incorporate MCTS as an inference-time decision mechanism. At each decision point, multiple hypotheses are generated and evaluated using the Upper Confidence Bound applied to Trees (UCT) algorithm (Equation 1[5], enabling the framework to balance exploration and exploitation. This search-guided process allows DrugMCTS to iteratively refine its predictions and autonomously filter low-quality or noisy information without requiring domain-specific fine-tuning.
By integrating structured data retrieval, expert-agent collaboration, and search-based decision-making, DrugMCTS offers a scalable and interpretable approach for scientific discovery in dynamic domains such as drug-target interaction.
@@ 2.1 Overview of DrugMCTS
Figure 3: Illustration of a single rollout in the DrugMCTS framework. The figure demonstrates how agents sequentially collaborate during one inference trajectory guided by MCTS.
2.2 MCTS for Drug Repositioning
Given that MCTS is a well-established algorithm widely adopted and thoroughly explained in various studies [15], we provide only an overview here. As shown in Figure 4 and Appendix Algorithm 2, the MCTS algorithm consists of four main steps: selection, expansion, simulation, and backpropagation.
@@ 2.3 Action Space
$A_{1}$ Retrieval Action. Upon receiving the query molecule, the Retrieval agent queries databases to identify molecules structurally similar to the query molecule. We employ two similarity metrics: the Tanimoto coefficient [1] and the cosine similarity based on the last hidden state computed by ChemBERTa. For each metric, we retrieve the top-10 most similar molecules, merge the results, and remove duplicates to form the initial candidate molecules $M_{cm}$. The proteins can interact with $M_{cm}$ are also retrieved, forming candidate proteins $P_{cp}$. These molecules serve as inputs to the root node of the MCTS.
@@
$A_{2}$ Molecule Analysis Action. Understanding molecular properties is crucial to predict molecular-target interactions [38]. While general-purpose LLMs can perform some analysis using SMILES representations, their interpretations are often incomplete and prone to errors, especially when experimental data, such as hydrophobicity, is required.
@@
$A_{3}$ Molecule Selection Action. Research has shown that the quality of retrieved information significantly impacts the accuracy of model-generated answers [22]. Excessive irrelevant information can negatively affect the model’s performance [18]. Although the retrieved molecules share structural similarities with the query molecule, they may not necessarily provide useful insights for drug repositioning tasks [14]. Thus, the Molecule Selection (MS) agent filters the molecule pool, based on structural similarity to the query molecule, pharmacophore integrity, and drug-like essentials, to generate reference molecules $M_{rm}$. The reference proteins $P_{rp}$ are obtained by selecting the proteins that can interact with $M_{rm}$ from $P_{cp}$. To ensure that these reference molecules are thoroughly characterized, this action also invokes the $A_{2}$ Molecule Analysis Action. Specifically, the MA agent is called to retrieve the structural and physicochemical properties ($C_{c,s}$ $C_{c,phy}$) of $M_{cm}$ but without generating reports.
@@ 2.4 Reward Calculation
Molecular Docking Methods. Molecular docking methods, such as AutoDock Vina [8], are widely used but suffer from notable limitations. First, they require accurate three-dimensional structural information of the target proteins, which is often unavailable or unreliable for many biologically relevant molecules. Second, these methods are computationally intensive, limiting their scalability in large-scale drug repositioning tasks. Given these limitations, there has been a growing interest in developing more efficient alternatives.
Deep Learning-Based Methods. Deep learning models have emerged as a promising solution due to their superior computational efficiency [32]. However, their performance is highly dependent on the training dataset. When predicting samples that deviate significantly from the training data, the accuracy of these models can drop by over 20% [31], limiting their applicability across diverse scenarios.
@@ 3.1 Datasets and Metrics
We utilized two datasets, DrugBank [12] and KIBA [25], which were processed to include a total of 788 entries from DrugBank and 626 entries from KIBA. Each entry consists of a molecule as input and its corresponding interacting proteins as output. The number of ground truth interactions per entry varies. To evaluate model performance, we used recall, defined as the ratio of correctly predicted proteins to the total number of ground truth proteins. The detailed information of data processing pipeline can be found in Appendix Section 0.A.1
$\text{Recall}=\frac{|\{\text{proteins predicted correctly}\}|}{|\{\text{all ground truth proteins}\}|}$ (7)
@@ 3.3 Baselines
General Models with RAG. (GM + RAG) Many existing studies on drug repositioning using LLMs either employ divergent methodological formulations or are not open-sourced, which complicates comparative evaluation. Our approach adopts the framework of the DrugReAlign [29] method, with modifications and enhancements tailored to our specific formulation. To ensure equitable comparison, we incorporate not only protein pocket information but also incorporate $C_{q,s}$ and $C_{q,phy}$ as additional features. The models used in this group remained GPT-4o-mini and Deepseek-R1.
@@ 3.4 Results
The experimental results (Table 1) indicate that general-purpose LLMs (GPT-4o-mini and Deepseek-R1) achieved relatively low recall scores of only 12.59%–16.19% on the DrugBank dataset when operating in a zero-shot setting. However, when incorporating molecular structural features and chemical properties via RAG-based prompting, model performance decreased, with GPT-4o-mini dropping to 15.19% and Deepseek-R1 significantly falling to 12.59%. This suggests that the inclusion of potentially irrelevant or misleading information through retrieval can negatively impact the reasoning capabilities of general-purpose LLMs.
@@
These results strongly underscore the superiority of dynamic decision-making mechanisms, such as those used in DrugMCTS, over traditional static prediction methods like deep learning models. They also highlight the limitations of general-purpose large models in zero-shot settings for drug repositioning tasks, especially when retrieval-augmented prompting introduces noise or irrelevant context. This further emphasizes the importance of structured reasoning, domain-specific knowledge integration, and information filtering in such applications.
Furthermore, to demonstrate the interpretability and transparency of our approach, we provide a detailed case study in the Appendix Section 0.A.2. It illustrates how the model predicts an interaction between Equol and the CXC chemokine receptor 3, including the complete step-by-step reasoning process during a specific MCTS rollout.